### v2

# EvolveNav: Self-Improving Embodied Reasoning for LLM-Based Vision-Language Navigation

Abstract
Building Vision-Language Navigation (VLN) agents which can navigate following natural language instructions is a long-standing goal in human-robot interaction applications. Recent studies have revealed the potential of training open-source Large Language Models (LLMs) to unleash LLMs’ reasoning ability for improving navigation, and simultaneously mitigate the domain gap between LLMs’ training corpus and the VLN task. However, these approaches primarily adopt direct input-output mapping paradigms, causing the mapping learning difficult and the navigational decisions unexplainable. Chain-of-Thought (CoT) training is a promising way to improve both navigational decision accuracy and interpretability, while the complexity of the navigation task makes the perfect CoT labels unavailable and may lead to overfitting through pure CoT supervised fine-tuning. In this paper, we propose a novel sElf-improving embodied reasoning framework for boosting LLM-based vision-language Navigation, dubbed EvolveNav. Our EvolveNav consists of two stages: (1) Formalized CoT Supervised Fine-Tuning, where we train the model with formalized CoT labels to both activate the model’s navigational reasoning capabilities and increase the reasoning speed; (2) Self-Reflective Post-Training, where the model is iteratively trained with its own reasoning outputs as self-enriched CoT labels to enhance the supervision diversity. A self-reflective auxiliary task is also introduced to encourage learning correct reasoning patterns by contrasting with wrong ones. Experimental results on the popular VLN benchmarks demonstrate the superiority of EvolveNav over previous LLM-based VLN approaches. Code is available at: https://github.com/expectorlin/EvolveNav.
1 Introduction
Vision-Language Navigation (VLN) has received significant research interest within the Embodied AI community, due to its practicality and flexibility in enabling human-robot interaction in real-world robotic applications. In VLN tasks, an embodied agent needs to follow natural language instructions to navigate through complex visual environments to reach the target position. Early works improve VLN performance by designing dedicated model architectures wang2019reinforced ; ma2019self ; deng2020evolving ; qi2020Object , introducing powerful learning paradigms zhu2020vision ; li2019robust ; huang2019transferable , and developing useful data augmentation techniques tan2019learning ; fried2018speaker ; liu2021vision ; Fu2019CounterfactualVN . Subsequently, pretraining-based VLN approaches have been widely proposed to improve the cross-modal alignment ability and decision accuracy of navigation agents hong2021vln ; Chen2021HistoryAM ; Qi2021TheRT ; Chen2022ThinkGA ; Qiao2022HOPHA . Nevertheless, constrained by the limited scale of pretraining and VLN in-domain data, these approaches cannot learn navigational reasoning knowledge sufficiently, and therefore still struggle to handle various unseen navigation scenarios.
With the rapid progress of large language models (LLMs) brown2020language ; touvron2023llama ; touvron2023llama2 , emerging works have introduced LLMs to address embodied tasks considering LLMs’ rich real-world common sense and powerful reasoning ability Ahn2022DoAI ; Huang2022InnerME ; driess2023palme . Some recent works have attempted to build LLM-based VLN models in a zero-shot or trainable manner  chen2024mapgpt ; zhou2023navgpt ; long2023discuss ; hao2020towards ; zhang2024navid . The zero-shot approaches, such as NavGPT zhou2023navgpt and MapGPT chen2024mapgpt , resort to closed-source LLMs OpenAI_2023 to generate navigational reasoning and decision for different navigation timesteps. To alleviate the high cost of frequently querying closed-source LLMs for sequential decision making, the trainable methods collect in-domain data to train open-source LLMs vicuna to build the navigation agent. However, they typically map directly from navigational inputs to decisions without explicit intermediate reasoning steps, leading to decision uninterpretability and may also limit performance (see Figure 1(a)).
Figure 1: Comparison of different chain-of-thought (CoT) training paradigms. (a) Direct Mapping Learning maps the navigation inputs to actions straightforwardly. (b) Formalized CoT Learning and (c) Free-form CoT Learning generate formalized and free-form reasoning, respectively, under the training with fixed CoT labels. (d) Different from the above paradigms, our Self-Improving CoT Learning framework utilizes the model’s own reasoning outputs as self-enriched CoT labels and learn the reasoning in a self-reflective way during CoT training to fulfill generalizable and adaptable reasoning. Red and green fonts represent wrong and correct reasoning outputs, respectively. R+ and R- represent positive and negative reasoning samples, respectively.
Recent studies have revealed the effectiveness of embodied chain-of-thoughts (CoT) training in enhancing both the decision accuracy and interpretability for embodied tasks zawalski2024robotic ; zhao2025cot ; lin2024navcot . Most of these approaches employ the supervised fine-tuning paradigm for embodied CoT training. However, introducing CoT supervised fine-tuning for training VLN models is challenging. Firstly, due to the complexity and uncertainty of the navigation task, there can be multiple cues for deciding the correct navigation action, i.e., there can be multiple “correct” CoT supervisions to guide navigation for a specific timestep. This leads to a hard collection process of perfect navigation CoT labels. Secondly, pure CoT supervised fine-tuning using fixed CoT labels may cause overfitting to certain reasoning patterns and harm the generalization to diverse unseen scenarios.
In this paper, we propose a sElf-improving embodied reasoning paradigm for enhancing LLM-based vision-language Navigation, called EvolveNav, to fulfill generalizable and adaptable navigational reasoning under different scenes. Our EvolveNav comprises two stages: 1) Formalized CoT Supervised Fine-Tuning and 2) Self-Reflective Post-Training. The Stage 1 training of Formalized CoT Supervised Fine-Tuning aims to first activate the model’s potential reasoning capabilities, where we ask the model to produce explicit chain-of-thought navigational reasoning by predicting the landmarks needed to locate with the corresponding direction for deciding the navigation actions. To alleviate redundant reasoning and increase the inference speed, we conduct the CoT supervised fine-tuning using formalized CoT labels, which are collected by filling the landmark and direction information into concise label templates. To mitigate the overfitting to pre-constructed CoT labels and enable self-improving reasoning for enhancing generalization, we further conduct Self-Reflective Post-Training for Stage 2 training. Concretely, we design a self-enriched CoT label enhancement strategy, where we train the model with its iteratively produced correct reasoning outputs to enrich the CoT supervision. We also introduce a self-reflective auxiliary task, where the model needs to discriminate between positive and negative navigational reasoning to learn correct reasoning patterns better. As shown in Figure 1, in contrast to CoT supervised fine-tuning using fixed CoT labels (i.e., Figure 1(b) Formalized CoT Learning and (c) Free-form CoT Learning), our EvolveNav (Figure 1(d)) can generate embodied CoT in a self-refining manner during training to mitigate the overfitting. Additionally, through training with formalized CoT labels, our EvolveNav can significantly reduce uninformative navigational reasoning to promote the reasoning speed compared with using free-form CoT labels for training (as in Figure 1 (c)).
We conduct experiments on two public VLN benchmarks, R2R anderson2018vision and CVDN thomason2019vision . Experimental results show that EvolveNav substantially outperforms previous LLM-based VLN approaches on unseen scenarios, demonstrating the effectiveness of our self-improving embodied reasoning paradigm in promoting navigation decision accuracy and generalization. We carefully conduct ablation experiments to explore how to design streamlined CoTs that can provide interpretability while boosting navigation performance. Visualization also insightfully reveals the reasonability of our design for CoT labels in enhancing decision interpretability and improving navigational reasoning.
To summarize, the main contributions of this paper are:
  • We propose a novel self-improving embodied reasoning paradigm, EvolveNav, for enhancing the LLM-based vision-and-language navigation, which fulfills generalizable and adaptable navigational reasoning under unseen scenes.
  • We construct formalized CoT labels for conducting supervised fine-tuning, which effectively activates the agent’s navigational reasoning ability and promotes the reasoning speed. We introduce a self-enriched CoT label enhancement strategy and a self-reflective auxiliary task to enable learning correct reasoning patterns in a self-refining manner to mitigate overfitting.
  • Experimental results show the superiority of EvolveNav over other LLM-based approaches on two standard VLN benchmarks R2R and CVDN. Our EvolveNav can improve the generalization of both navigational reasoning and decision-making, providing meaningful insights for designing further embodied reasoning paradigms.
2 Related Work
Vision-Language Navigation (VLN). Vision-Language Navigation (VLN) has attracted intensive research interest in recent years. Various VLN benchmarks have been proposed to evaluate agents’ ability for navigational reasoning and instruction following anderson2018vision ; jain2019stay ; ku2020room ; thomason2019vision ; qi2020reverie ; zhu2021soon . Previous approaches employ non-pretraining-based wang2019reinforced ; ma2019self ; deng2020evolving ; zhu2020vision ; tan2019learning ; fried2018speaker or pretraining-based paradigms hong2021vln ; Chen2021HistoryAM ; Chen2022ThinkGA ; Qiao2022HOPHA ; Guhur2021AirbertIP ; an2022bevbert ; wang2023scaling for tackling the above VLN tasks. However, these approaches cannot generalize well to diverse unseen scenarios that require rich real-world commonsense, and the navigation decisions also lack explainability. Some recent works have introduced LLMs to assist the VLN task, by either eliciting the useful navigation knowledge stored in LLMs lin2024correctable ; zhou2024navgpt ; qiao2023march or employing the LLM as the navigation backbone for action decision zhou2023navgpt ; zheng2024towards ; zhang2024navid ; lin2024navcot ; chen2024mapgpt . Our work lies in the latter. Different from previous approaches, we propose a new VLN framework in this work, where the LLM-based navigation backbone iteratively generates intermediate reasoning steps in a self-improving manner during training to guide navigational decisions. As a result, both the reasoning ability and decision interpretability of the navigation model can be significantly enhanced.
LLMs as Embodied Agents. Recent research have revealed the giant potential of utilizing Large Language Models (LLMs) as embodied agents to complete the robotic navigation and manipulation tasks, benefiting from the outstanding ability of planning, reasoning, and reflection of LLMs song2023llmplanner ; Ahn2022DoAI ; Huang2022InnerME ; Yao2022ReActSR ; shahlm ; schumann-2023-velma ; wang2023voyager ; huang2023voxposer ; rajvanshi2024saynav . For example, LM-Nav shahlm introduces the LLM to parse the long navigation instruction into sequential landmarks for facilitating the navigational planning. Voxposer huang2023voxposer introduces LLMs for code writing and combines them with the Vision-Language models (VLMs) to compose 3D value maps for robotic manipulation.
There are typically two branches of works where the LLMs act as embodied agents for tackling the VLN task. In the first branch, closed-source LLMs like GPT-4 openai are queried in a zero-shot manner to decide the action sequentially zhou2023navgpt ; long2023discuss ; chen2024mapgpt ; chen2024affordances . For example, NavGPT zhou2023navgpt transforms visual observations into textual formats and feed them to LLMs for generating action predictions. The second branch finetunes open-source LLMs with in-domain VLN datasets, which alleviates the LLM’s query cost as well as mitigates the gap between LLM’s training corpus and VLN tasks zheng2024towards ; lin2024navcot ; zhang2024navid ; han2024roomtour3d . For example, Navid zhang2024navid constructs a video-based navigational vision-language model and train it using navigation samples collected from continuous R2R datasets. However, most of them map navigation inputs to action decisions without the reasoning output. In contrast, we train the open-source LLM to generate self-improving embodied reasoning explicitly to improve action decision accuracy, which enhances the decision interpretability as well as mitigates overfitting to training reasoning labels.
Embodied Chain-of-thoughts Training. Chain-of-thoughts (CoT) reasoning wei2022chain has been a widely utilized technique in Large Language Models (LLMs) and Vision-Language Models (VLMs). By generating the intermediate reasoning steps rather than directly predicting the answer, CoT can promote the answer accuracy for various tasks such as mathematical reasoning, commonsense reasoning, code generation, etc openai ; OpenAI_2023 ; liu2024deepseek . Inspired by this, some recent works have trained LLMs/VLMs to generate reasonable CoT for improving the action decision accuracy in embodied tasks zawalski2024robotic ; lin2024navcot ; mu2023embodiedgpt ; zhao2025cot ; liu2025spatialcot ; zhou2024navgpt . ECoT zawalski2024robotic trains vision-language-action models to generate embodied reasoning including object bounding boxes and end effector positions to encourage better adaptation for robotic manipulation tasks. NavGPT-2 zhou2024navgpt resorts to the GPT-4V model to collect free-form step-wise CoT reasoning data to improve the navigational reasoning ability of the VLN-specialized models Chen2022ThinkGA . CoT-VLA zhao2025cot incorporates explicit visual chain-of-thought reasoning into vision-language-action models by predicting future image frames before generating the action sequence. In our work, we propose a novel self-improving embodied CoT training paradigm for boosting the VLN, which effectively improves the reasoning ability and decision accuracy of the navigation model in various unseen scenarios. Moreover, we construct formalized CoT supervision, which can reduce redundant reasoning information and simultaneously promote the model’s inference speed.
Figure 2: Overview of our EvolveNav. EvolveNav introduces a two-stage training framework for fulfilling self-improving embodied reasoning. In Stage 1 Formalized CoT Supervised Fine-Tuning, the navigation agent is trained using pre-constructed formalized CoT labels to generate navigational reasoning by predicting the landmark needed to locate with the corresponding direction. In Stage 2 Self-Reflective Post-Training, the agent’s own reasoning outputs are introduced as the self-enriched CoT labels to enhance the supervision diversity. A self-reflective auxiliary task is also designed to guide the navigation agent to discriminate between correct and wrong reasoning outputs.
3 Method
In this section, we first introduce the problem definition of the VLN task (Sec. 3.1). Then, we present the model architecture of the LLM-based navigation agent in our EvolveNav (Sec. 3.2). Finally, we describe our proposed self-improving embodied reasoning framework in detail (Sec. 3.3). The overview of our EvolveNav is presented in Figure 2. Specifically, our EvolveNav consists of two stages: 1) Formalized CoT Supervised Fine-Tuning (Sec. 3.3.1), where the LLM-based VLN model is initially trained using formalized CoT labels with supervised fine-tuning to activate the model’s navigational reasoning ability and simultaneously promote the reasoning speed; 2) Self-Reflective Post-Training (Sec. 3.3.2), where the model is further trained with its own reasoning outputs as self-enriched CoT labels to increase supervision diversity, accompanied by a self-reflective auxiliary task to encourage better learning of accurate navigational reasoning patterns by discriminating from incorrect ones.
3.1 Problem Setup
In the VLN task, an agent is given a navigation instruction $I$ in the form of a declarative sentence or a dialogue and is required to navigate from a start position to the target position. At timestep $t$, the agent receives a panoramic observation $O_{t}$ containing $K$ single-view observations $O_{t,k}$, i.e., $O_{t}=\{O_{t,k}\}_{k=1}^{K}$. There are $N$ navigable views among $K$ views. The navigable views and the stop action form the action space, from which the agent chooses one as the action prediction $a_{t}$. Actions before step $t$ are treated as the navigation history.
3.2 Model Architecture
We build the LLM-based navigation agent that can simultaneously produce the navigational chain-of-thought reasoning and action prediction, modifying from a recent LLM-based VLN work, NaviLLM zheng2024towards . Concretely, the navigation agent consists of a scene encoder $F_{v}$, an LLM backbone $F_{\mathrm{LLM}}$, and an action prediction head $F_{\mathrm{action}}$. At each timestep $t$, the agent receives the navigation instruction $I$, panoramic observation $O_{t}$, and navigation history features $H_{t}=\{h_{0},...,h_{t-1}\}$. The scene encoder transforms $N$ navigable panoramic views $\{O_{t,n}\}_{n=1}^{N}$ into visual representations $\{V_{t,n}\}_{n=1}^{N}$:
$\{V_{t,n}\}_{n=1}^{N}$ $=F_{v}(\{O_{t,n}\}_{n=1}^{N}).$ (1)
The navigation prompt $P$ is then constructed by integrating the tokenized instruction, the visual representations $\{V_{t,n}\}_{n=1}^{N}$, and navigation history features $H_{t}$. As shown in Figure 2, special tokens <hist> and <cand> are introduced as placeholder tokens, where we insert the features $H_{t}$ and $\{V_{t,n}\}_{n=1}^{N}$, respectively. In contrast to NaviLLM zheng2024towards that directly maps the navigational inputs to action decision, in our EvolveNav, we construct the following output hint in the prompt $P$ to guide the navigation agent to generate both the action decision and explicit navigational reasoning: “-Action Decision: <cls>. -Navigational Reasoning: ”. The <cls> token is also a special token for facilitating subsequent action predictions. The navigation prompt $P$ is fed into the LLM backbone $F_{\mathrm{LLM}}$ to obtain the feature ${f}^{cls}_{t}$ of the <cls> token and the chain-of-thought (CoT) reasoning $\mathrm{CoT}$:
${f}^{cls}_{t},\mathrm{CoT}$ $=F_{\mathrm{LLM}}(P).$ (2)
Under the guidance of CoT reasoning, the ${f}^{cls}_{t}$ is sent to the action prediction head $F_{\mathrm{action}}$ for generating the action prediction $a_{t}$:
${a}_{t}$ $=F_{\mathrm{action}}({f}^{cls}_{t}).$ (3)
3.3 Self-Improving Embodied Reasoning
3.3.1 Stage 1: Formalized CoT Supervised Fine-Tuning
Formalized CoT Labels Collection. When facing a given human instruction, the navigation agent usually needs to sequentially reason about the direction or the landmark it should move to in its current visual observation to reach the target position. Therefore, in our EvolveNav, we train the LLM-based VLN model to generate the chain-of-thought (CoT) reasoning about the landmark with the corresponding direction at different navigation timesteps, like the following format: “I should go to an observation with [landmark] to the [direction] of me”.
To encourage the navigation agent to choose the ground-truth action $a^{*}_{t}$ (paired observation is denoted as $O_{t}^{*}$) at different timesteps $t$ through the guidance of CoT reasoning during training, we obtain the corresponding landmarks $L$ and direction $D$ of $O_{t}^{*}$ to construct the formalized CoT labels, which is described as follows. Denote the observation $O_{t}^{*}$ as $O_{t}^{*}=\{B_{t},A_{t}=\{\psi_{t},\theta_{t}\}\}$, where $B_{t}$ is the RGB image of $O_{t}^{*}$, $A_{t}$ represent the direction information containing heading $\psi_{t}$ and elevation $\theta_{t}$. For the image $B_{t}$, we first employ a powerful image captioning model huang2023tag2text $F_{\mathrm{cap}}$ to obtain object and scene context $C_{t}$:
$C_{t}=F_{\mathrm{cap}}(B_{t}).$ (4)
Then, we leverage the NLP tool Spacy spacy to extract the landmarks list $L$ from $C_{t}$. In contrast to directly using object recognition models which may detect multiple redundant objects, extracting landmarks from the image captions can better retain salient landmarks. As a result, the generated CoT reasoning of the navigation agent can effectively help it locate important landmarks mentioned in the human instruction, since humans also tend to focus on salient landmarks when giving navigation instructions. We follow lin2024navcot to map the direction information $A_{t}$ of the observation $O_{t}^{*}$ to textual represented direction $D$. With the landmarks list $L$ and direction $D$, we construct CoT labels $\mathrm{\mathrm{CoT}^{*}}$ by filling the following label template: I should go to an observation with [$L$] to the [$D$] of me. Through using our formalized CoT labels for training, the generation of redundant reasoning information can be effectively alleviated to promote the reasoning speed.
Supervised Fine-Tuning with Formalized CoT Labels. To activate the potential navigational reasoning ability of the LLM agent to adapt to the VLN task, we introduce the supervised-finetuning (SFT) paradigm in Stage 1 for conducting CoT training with our pre-constructed formalized CoT labels. Denote the navigation data sample at each timestep $t$ as $(P,\mathrm{CoT^{*}})$ (we omit the subscript $t$ for simplicity), where $P$ and $\mathrm{CoT^{*}}$ are the navigation prompt and CoT label, respectively. The training objective $\mathcal{L}_{\mathrm{SFT}}$ maximizes the likelihood of generating $\mathrm{CoT}^{*}$ given $P$ auto-regressively:
$\mathcal{L}_{\mathrm{SFT}}=-\mathbb{E}_{(P,\mathrm{CoT}^{*})\sim\mathcal{D}}\sum_{s=1}^{S}\mathrm{log}F_{LLM}(\mathrm{CoT}^{*}_{s}|P,\mathrm{CoT}^{*}_{<s}),$ (5)
where $\mathcal{D}$ represents the navigation dataset. Denote the navigation action prediction training objective as $\mathcal{L}_{\mathrm{action}}$, the total training objective $\mathcal{L}_{\mathrm{Stage1}}$ of Stage 1 is calculated as follows:
$\mathcal{L}_{\mathrm{Stage1}}=\mathcal{L}_{\mathrm{action}}+\lambda\mathcal{L}_{\mathrm{SFT}},$ (6)
where $\lambda$ represents the loss balance factor. We follow zheng2024towards to calculate the action prediction training objective $\mathcal{L}_{\mathrm{action}}$.
3.3.2 Stage 2: Self-Reflective Post-Training
Although CoT supervised fine-tuning can explicitly guide the navigation agent to produce navigational reasoning for assisting the action decision, due to the uncertainty and complexity of the navigation task, using fixed labels may lead to overfitting to training CoT label distributions and therefore harm the generalization to unseen scenarios. Moreover, the inherent noise in the image captioning model huang2023tag2text for landmark detection may also limit the accuracy of formalized CoT labels collected in Stage 1. Therefore, after the Stage 1 training of Formalized CoT Supervised Fine-Tuning, we introduce Self-Reflective Post-Training to further encourage the navigation agent to learn correct reasoning patterns for improving generalization.
Self-Enriched CoT Label Enhancement. To mitigate the overfitting to fixed CoT labels during training, we utilize the model’s self-generated reasoning outputs as self-enriched CoT labels under the guidance of the model’s action decision. At timestep $t$, denote the model’s reasoning output as $R_{t}$, the original formalized CoT label as $\mathrm{CoT}_{t}^{*}$, the model’s action decision as $a_{t}$, and the ground-truth action as $a_{t}^{*}$. When the action decision $a_{t}$ generated by the navigation agent matches the ground-truth action $a_{t}^{*}$, we choose the agent’s own reasoning output $R_{t}$ as the new CoT label. Such self-enriched CoT labels can effectively enhance the supervision diversity in a decision-oriented manner. Concretely, we obtain the updated CoT label $\tilde{\mathrm{CoT}}_{t}^{*}$ at timestep $t$ through the following rules:
$\tilde{\mathrm{CoT}}_{t}^{*}=\begin{cases}R_{t},\quad\quad\text{if}\quad a_{t}=a_{t}^{*}\\ \mathrm{CoT}_{t}^{*},\quad\text{otherwise}\end{cases}$ (7)
Self-Reflective Auxiliary Task. To further make the model aware of correct and wrong reasoning, which can help the model better learn correct reasoning patterns, we additionally introduce a self-reflective auxiliary task, where we ask the model to discriminate which reasoning output from the given reasoning is right. Specifically, we collect positive and negative reasoning samples $R^{+}$ and $R^{-}$ during training for conducting the self-reflective auxiliary task. We utilize the above mentioned CoT label $\tilde{\mathrm{CoT}}_{t}^{*}$ as the positive reasoning sample $R^{+}$. To obtain the negative reasoning sample $R^{-}$, we randomly select the candidate observation (action) $O_{t,j}$ ($1<j<N$, $N$ is the number of navigable views) different from the ground-truth one. Then we extract the landmark and direction for $O_{t,j}$ to fill in the CoT label template (see Sec. 3.3.1). As shown in Figure 2, we construct the self-reflective task prompt $P_{\mathrm{sr}}$ as “Choose the correct one from the given two navigational reasoning outputs. Output 1: [$R^{1}$]. Output 2: [$R^{2}$]. Selection: ”, where we randomly insert the positive reasoning sample $R^{+}$ and negative reasoning sample $R^{-}$ to the positions of $R^{1}$ and $R^{2}$. We collect the ground-truth output $R_{\mathrm{sr}}^{*}$ with the form like “Output 2.”. We also utilize the auto-regressive training objective like Eq. 5 to calculate the loss $\mathcal{L}_{\mathrm{sr}}$ based on ($P_{\mathrm{sr}}$, $R_{\mathrm{sr}}^{*}$) pairs for the self-reflective auxiliary task.
With the self-reflective loss $\mathcal{L}_{\mathrm{sr}}$, the total training objective $\mathcal{L}_{\mathrm{Stage2}}$ for Stage 2 is obtained by:
$\mathcal{L}_{\mathrm{Stage2}}=\mathcal{L}_{\mathrm{action}}+\lambda_{1}\mathcal{L}_{\mathrm{SFT}}+\lambda_{2}\mathcal{L}_{\mathrm{sr}},$ (8)
where both $\lambda_{1}$ and $\lambda_{2}$ are the loss coefficients. Through the self-enriched CoT label enhancement and self-reflective auxiliary task, the navigation agent learns to generate correct embodied reasoning in a self-refining manner to enable adaptable reasoning in different scenarios.
4 Experiment
4.1 Experimental Setup
Datasets. We test our EvolveNav on two popular VLN benchmarks, i.e., R2R anderson2018vision and CVDN thomason2019vision . R2R is built on 90 real-world indoor simulation environments containing 7,189 trajectories, each corresponding to three fine-grained instructions. CVDN contains 2,050 human-human navigation dialogs and over 7k trajectories in 83 MatterPort houses.
Evaluation Metrics. We utilize the following standard metrics for evaluation: 1) Trajectory Length (TL): the average length of the agent’s navigated path, 2) Navigation Error (NE): the average distance between the agent’s destination and the goal viewpoint, 3) Success Rate (SR): the ratio of success, where the agent stops within three meters of the target point, 4) Success rate weighted by Path Length (SPLanderson2018vision : success rate normalized by the ratio between the length of the shortest path and the predicted path, 5) Oracle Success Rate (OSR): the ratio of containing a viewpoint along the path where the target position is visible, 6) Goal Progress (GP), the progress in meters towards the goal.
Table 1: Comparison with state-of-the-art methods on the R2R dataset. * denotes training with R2R in-domain data. IL means the imitation learning setting.
Method Val Unseen Test Unseen
[…]
Cross-Modal Backbone:
PREVALENT hao2020towards 10.19 4.71 58 53 10.51 5.30 54 51
HOP qiao2022hop 12.27 3.80 64 57 12.68 3.83 64 59
HAMT Chen2021HistoryAM 11.46 2.29 66 61 12.27 3.93 65 60
VLN-BERT hong2021vln 12.01 3.93 63 57 12.35 4.09 63 57
DUET Chen2022ThinkGA 13.94 3.31 72 60 14.73 3.65 69 59
Meta-Explore Hwang2023MetaExploreEH 13.09 3.22 72 62 14.25 3.57 71 61
VLN-SIG li2023improving - - 72 62 - - 72 60
VLN-PETL qiao2023vln 11.52 3.53 65 60 12.30 4.10 63 58
NavGPT2 zhou2024navgpt 13.25 3.18 71 60 - - - -
LLM-based Backbone:
NavGPT zhou2023navgpt 11.45 6.46 34 29 - - - -
DiscussNav long2023discuss 9.69 5.32 43 40 - - - -
MapGPT chen2024mapgpt - 5.63 34 29 - - - -
NavCoT lin2024navcot 9.95 6.26 40 37 - - - -
NaviLLM zheng2024towards 12.81 3.51 67 59 13.21 3.71 68 60
NaviLLM* (IL) 9.99 6.04 46.90 43.78 10.03 6.12 46 43
EvolveNav (IL) 9.79 5.52 51.15 48.27 9.94 5.92 47 45
NaviLLM* 13.43 3.27 70.11 60.25 13.68 3.37 70 61
EvolveNav 12.07 3.15 71.17 63.48 12.06 3.22 71 63
Implementation Details. We fine-tune the LLM with full-parameter and LoRA settings for Stage 1 and 2, respectively. To simplify the implementation, we do not introduce the pretraining phase like zheng2024towards for both our EvolveNav and the compared baseline (denoted as NaviLLM* in Table 1 &  2). The training for Stage 1 is conducted on 8 Nvidia A100 GPUs and the training for Stage 2 is performed on 4 Nvidia A100 GPUs. Empirically, we set the loss coefficients $\lambda$, $\lambda_{1}$, and $\lambda_{2}$ as 1, 1, and 0.2, respectively. The hyperparameters such as the learning rate, optimizer, and the batch size are kept the same as zheng2024towards .More implementation details are presented in the supplementary material.
4.2 Comparison with Existing Methods
Table 1 shows the comparison among different approaches on R2R. We compare EvolveNav with the baseline method NaviLLM zheng2024towards using the R2R in-domain training data (denoted as NaviLLM*) under two settings, i.e., imitation learning (IL) and Dagger training zheng2024towards ; Chen2022ThinkGA . The comparison shows the consistent improvement of our approach over NaviLLM* on both Val Unseen and Test Unseen under different training settings. For example, the performance gain in SPL of EvolveNav on Val Unseen is $\sim$4.5and $\sim$3.2under IL and dagger training settings compared to the baseline, respectively. Note that our EvolveNav achieves the state-of-the-art performance on R2R among the approaches with LLM-based backbone. These results demonstrate the effectiveness of our self-improving embodied reasoning framework in fulfilling generalizable navigational reasoning. We do not consider the comparison with models augmented by new environments (e.g., ScaleVLN wang2023scaling ) for fairness.
Table 2: Comparison with SoTA methods on CVDN. We utilize the Goal Progress (GP) (m) as the evaluation metric. * denotes training with CVDN in-domain data.
Method Val-Unseen Test
Cross-Modal Backbone:
Seq2Seq thomason2019vision 2.10 2.35
PREVALENT hao2020towards 3.15 2.44
HOP qiao2022hop 4.41 3.31
MT-RCM wang2020environment 4.36 -
MT-RCM+Env wang2020environment 4.65 3.91
HAMT Chen2021HistoryAM 5.13 5.58
VLN-SIG li2023improving 5.52 5.83
VLN-PETL qiao2023vln 5.69 6.13
LLM-based Backbone:
NaviLLM zheng2024towards 6.16 7.90
NaviLLM* 5.53 6.80
EvolveNav (ours) 6.21 7.07
In Table 2, we compare our EvolveNav with different VLN approaches on CVDN. As shown in Table 2, EvolveNav outperforms the baseline NaviLLM* zheng2024towards on both Val Unseen and Test set on CVDN, demonstrating good generalization ability of our proposed self-improving embodied reasoning paradigm to different kinds of VLN tasks. For example, EvolveNav brings $\sim$0.7 performance gain over NaviLLM* in GP on Val Unseen. Moreover, Table 2 shows the superiority of EvolveNav over NaviLLM zheng2024towards , which conducts multi-task training on multiple VLN benchmarks to learn cross-task navigation knowledge, demonstrating the importance of explicit navigational reasoning learning compared to direct input-output mapping paradigms. Such results also reveal the effectiveness of our constructed CoT labels, which provide landmark and direction information for enabling the agent to learn useful navigational reasoning for decisions. The results on both R2R and CVDN benchmarks demonstrate the effectiveness of our EvolveNav in fulfilling generalizable navigational reasoning to improve navigation decision accuracy.
4.3 Ablation Study
Effect of different method components. Table 3 presents the ablation study on Val Unseen set on R2R, where we can find the effectiveness and reasonability of different method components in our EvolveNav. From Table 3, we can observe that through the Stage 1 training of formalized CoT supervised fine-tuning (SFT), the model’s navigational reasoning ability can be significantly enhanced to improve the navigation performance. The introduction of self-enriched CoT labels and the self-reflective auxiliary task can further bring performance gain in both SR and SPL metrics compared to pure CoT SFT. Our full model achieves the best results in SR and SPL among all variants, demonstrating the effectiveness of our two-phase self-improving embodied training framework.
Table 3: Ablation study of method components on Val Unseen set on R2R. We adopt the imitation learning (IL) setting for evaluation.
[…]
Baseline 46.90 54.63 43.78
   + Formalized CoT SFT 49.62 59.44 46.26
   + Self-Enriched CoT SFT 50.47 57.48 47.98
   + Self-reflective task 50.51 57.74 47.86
   + Full Model (ours) 51.15 59.18 48.27
Effect of constructed CoT labels. Table 4 compares the navigation performance under different kinds of CoT labels, where we can find the effectiveness of our introduced CoT labels by predicting landmark and direction information in a formalized way. To realize “Free-form CoT”, we introduce the free-form CoT labels collected in NavGPT-2 zhou2024navgpt to train the navigation agent. To realize “Direction & Landmark”, we obtain the best matched landmark in the instruction to each ground-truth observation through the CLIP model radford2021learning . The comparison between “Free-form CoT” and “Direction & Landmark (ours)” shows that our formalized CoT labels can effectively reduce redundant reasoning information to improve the navigational reasoning and decision accuracy. Moreover, the inference time to generate CoT at one timestep of “Free-form CoT” is $\sim$7.8s compared to $\sim$2.5s of “Direction & Landmark (ours)”, demonstrating that our method can significantly promote the reasoning speed ($\sim$×$\times$3 improvement), which is crucial for sequential decision making task like navigation. By comparing “Only Direction” and “Only Landmark” with “Direction & Landmark (ours)”, we can find that both landmark and direction information are important to guide navigation decisions, revealing the reasonability of our constructed CoT labels. The superiority of “Direction & Landmark (ours)” over “Direction & Landmark” demonstrates that our introduced CoT labels, which contain diverse landmarks, can potentially encourage the navigation agent to learn cross-modal alignment knowledge to accurately align the visual observation to the navigation instruction.
Table 4: Ablation study of CoT on Val Unseen set on R2R. denotes using the landmark mentioned in the instruction.
[…]
Baseline 70.11 79.00 60.25
Free-form CoT 69.39 80.44 60.95
Only Direction 67.05 76.49 57.32
Only Landmark 68.32 79.85 60.53
Direction & Landmark 69.77 79.25 60.77
Direction & Landmark (ours) 71.26 80.23 62.05
4.4 Visualization
Fig. 3 gives the action decision visualization comparison between NaviLLM zheng2024towards and our EvolveNav, where we can find that EvolveNav generates reasonable navigational reasoning about landmarks and directions to guide correct action decision making. For example, when $t=0$, from the observations, EvolveNav infers that an observation with door and hallway represents the exit from the bedroom while NaviLLM mistakenly selects an action that remains in the bedroom. Another example is in the hallway ($t=4$ for NaviLLM and $t=1$ for EvolveNav), EvolveNav generates reasoning consistent with the decision of entering the correct doorway on the left. However, NaviLLM chooses the wrong doorway to go back. These results highlight the effectiveness of our approach in improving navigational reasoning for accurate instruction understanding and action prediction.
Figure 3: Action decision visualization of NaviLLM zheng2024towards and our EvolveNav. We only extract two steps and display local candidate space for simplicity. Observations selected by EvolveNav (also are the ground-truth actions) and NaviLLM are annotated by green boxes and red boxes, respectively.
5 Conclusion
In this work, we propose EvolveNav, a novel self-improving embodied reasoning framework to fulfill generalizable and adaptable reasoning for enhancing LLM-based vision-and-language navigation. Through introducing the formalized CoT supervised fine-tuning and self-reflective post-training in the proposed framework, the agent’s navigational reasoning ability can be effectively enhanced while mitigating the overfitting to the training reasoning label distributions simultaneously to improve generalization. Experimental results on two popular benchmarks R2R and CVDN reveal the promising capability of our method in boosting the reasoning ability and decision accuracy for LLM-based navigation agents. We believe that our EvolveNav can provide meaningful references for designing self-improving embodied reasoning paradigms to benefit future LLM-assisted Embodied AI research.
Supplementary Material
Appendix A Overview
In these supplementary materials, we first analyze the broader impacts and limitations of our EvolveNav in Sec. B and Sec. C, respectively. Then, we describe more implementation details in Sec. D. Finally, we present more experimental results, including the landmark extraction visualization, self-enriched CoT label visualization, and loss & performance curves during Self-Reflective Post-Training, in Sec. E.
Appendix B Broader Impact
The scalable deployment of LLM-based navigation robots holds significant potential to enhance operational efficiency in public spaces, reduce human labor costs, and benefit diverse populations including the elderly and individuals with disabilities. Our proposed self-improving embodied reasoning framework can help improve both navigation success rate and decision interpretability of LLM-based navigation agents, therefore contributing to scalable application of LLM-based navigation robots which have high robustness and reliability. However, the inherent hallucination of LLMs may cause incorrect decisions in complex environments, leading to potential risks such as collisions. More efforts devoted to hallucination suppression of LLMs and safety promotion of navigation robots are encouraged for future research in the community.
Appendix C Limitations
(1) In the Stage 1 training of Formalized CoT Supervised Fine-tuning in our EvolveNav, we conduct full-parameter training for the open-source LLM, vicuna-7B [1], like the baseline method [66]. This process costs $\sim$1.5 day training on 8 NVIDIA A100 GPUs with 80G memory, which is relative resource-intensive compared to conventional VLN approaches [10, 9, 17]. In future work, we plan to introduce parameter-efficient training strategies or resort to more lightweight LLMs to reduce the computational resource consumption.
(2) During the formalized CoT label collection for supervised fine-tuning, we employ the open-vocabulary image captioning model [24] to obtain the caption for visual observations, followed by the NLP tool Spacy [20] to extract the landmark from the caption. This process is performed in an offline manner, which might introduce noise for the CoT label construction due to the possible inaccuracy of the image captioning model and the NLP tool. Future work includes utilizing more powerful open-vocabulary models or conducting human checking to promote the accuracy of the CoT labels.
Appendix D Implementation Details
We adopt Vicuna 7B [1] as our large language model (LLM) backbone, following [66]. For R2R [5], we also introduce the auxiliary tasks of trajectory summarization and embodied question answering in [66] during training. The maximum number of landmark extracted for constructing formalized CoT labels is set as 5 to retain useful information and simultaneously promote the reasoning speed. To mitigate the possible landmark missing during landmark detection due to the view deviation for a single candidate observation, we fuse the landmark information of candidate observations with the same direction to construct the CoT labels. We remove some uninformative landmarks such as “floor” and “ceiling” empirically during the landmark extraction. During Stage 1 training, we introduce the CoT supervised finetuning loss $\mathcal{L}_{\mathrm{SFT}}$ under a probability of 0.5 to mitigate the overfitting to the pre-constructed CoT labels. The maximum numbers of training steps for Stage 1 and 2 are set as 60000 and 9000 steps, respectively. Training for Stage 1 with full parameter lasts for $\sim$1.5 days with $\sim$73G GPU memory, and training for Stage 2 with LoRA lasts for $\sim$1 day with $\sim$30G GPU memory. The inference is conducted on 4 V100 GPUs with the evaluation batch size of 2. During inference, we set the LLM temperature to 0 and opt for a greedy strategy to generate action decisions.
Appendix E More Experimental Results
Landmark extraction visualization. We compare different methods for constructing the formalized CoT labels, to verify the reasonability of our landmark extraction strategy by combining image captioning model [24] with the NLP tool Spacy [20]. Fig. 4 presents the landmark extraction visualization comparison of three image captioning models, Tag2Text [24], BLIP-v2 [29], LLaVA 1.6 vicuna 13b [35], and an open-vocabulary object recognition model RAM [64]. Concretely, we provide four candidates in a navigational step for these methods and use Spacy [20] to extract the landmarks in the output caption (except for RAM of directly obtaining tagging). From Fig. 4, we can observe that Tag2Text can capture more informative landmarks while having less redundancy and illusion. For example, for “Cand 2”, Tag2Text correctly detects a kitchen and living room, a dining table, and chair, while BLIP-v2 only detects a large dining table. Although LLaVA 1.6 13b generates abundant captions, it brings noisy landmarks like the outdoor after noun phrases extraction. RAM also generates meaningless and non-existent landmarks, like floor and armchair. These results show that our combination of Tag2text model [24] and NLP tool for landmark extraction can effectively retain informative landmarks while reducing redundancy and illusion for constructing CoT labels, which can enable the agent to better learn cross-modal alignment between observations and instructions.
Figure 4: Landmark extraction visualization of different methods. We use green, red, and blue colors to distinguish informative, false, and uninformative landmarks, respectively.
Self-Enriched CoT label visualization. In Fig. 5, we present some visualization comparison examples of original CoT labels and self-enriched CoT labels. From Fig. 5, we can observe that our self-enriched CoT label enhancement strategy effectively increases the supervision diversity. For example, in Fig. 5(a) and (b), the self-enriched CoT labels from the model’s own reasoning outputs capture the landmark painting and chair which are not contained in the original CoT label, respectively. From Fig. 5(c) and (d), we can find that the LLM-based navigation agent can also recognize the attribute of the landmark, e.g., it recognizes wood paneling and arched window, which indicate the material and shape of the landmark, respectively. Such CoT labels can help navigation agent learn to follow more fine-grained instructions. Benefiting from our self-enriched CoT label, the LLM-based navigation agent can reduce the overfitting to the original CoT label distributions and learn more diverse cross-modal alignment knowledge, and therefore promote the generalization to unseen scenarios.
Figure 5: Self-enriched chain-of-thought (CoT) label visualization. Newly introduced landmarks in the self-enriched CoT label are highlighted in red fonts. GT action denotes the ground-truth action (observation). We omit the direction information in the CoT label.
(a) Loss variation
(b) Navigation Error (NE) variation
Figure 6: Loss and performance variation during Stage 2: Self-Reflective Post-Training. Low navigation error (NE) value indicates better results.
Loss & performance variation during Stage 2 training. Fig. 6 shows the loss and performance curves during Self-Reflective Post-Training (Stage 2). In Fig. 6, we can find that our self-reflective loss $\mathcal{L}_{\mathrm{sr}}$, the total training loss $\mathcal{L}_{\mathrm{Stage2}}$, and the navigation error (NE) have similar variation trends. Especially, both two loss curves and the NE curves achieve the lowest value around epoch 9. These results show the effect of our self-reflective auxiliary task during Self-Reflective Post-Training in improving the navigational reasoning and decision accuracy of the LLM-based navigation agent.